please show me how to solve this problem 2(12g^2-g)=15 using quadratic equation .
thank you vary much,
Summer Breeanne

Answer by stanbon(75887)

please show me how to solve this problem 2(12g^2-g)=15 using quadratic equation .
-------
24g^2 - 2g - 15 = 0
----------------------
g = [2 +- sqrt(4-4*24*-15)]/48
-----
g = [2 +- sqrt1444]/48
----
g = [2 +- 38]/48
---
g = (40/48) or x = -3/4
============================
Cheers,
Stan H.
=================